Atwich Corporation uses the weighted-average method in its process costing system. This month, the beginning inventory in the first processing department consisted of 600 units. The costs and percentage completion of these units in beginning inventory were: Cost percent complete Material Cost $10,700 70% Conversion Cost $13,900 50% A total of 5,100 units were started and 4,400 units were transferred to the second processing department during the month. The following costs were incurred in the first processing department during the month: Material cost $103,700 Conversion cost $ 180,900 The ending inventory was 75% complete with respect to materials and 10% complete with respect to conversion costs. Note: Your answers may differ from those offered below due to rounding error. In all cases, select the answer that is the closest to the answer you computed. To reduce rounding error, carry out all computations to at least three decimal places. The cost of ending work in process inventory in the first processing department according to the company's cost system is closest to: 1- Total cost to account for using weighted average units 2- Equivalent unit of production 3- Cost of units completed and transferred out 4- Cost per equivalent. Please show your calculation to the above items.
Answer #1

Calculation of Equivalent Annual Units

Using Weighted Average Method

Particulars Units % Materials EUP Materials % Conversion EUP Conversion
Units Completed and Transferred Out 4,400 100% 4,400 100% 4,400
Ending Work in Process Inventory 1,300 75% 975 10% 130
Total 5,700 5,375 4,530

EUP Stands for Equivalent Units of Production

EUP Materials = Units * % Materials

EUP Conversion = Units * % Conversion

Units in Ending Inventory = Units in Beginning Inventory + Unit Introduced during the Year - Units Transferred and Completed

Units in Ending Inventory = 600 + 5,100 - 4,400

Units in Ending Inventory = 1,300 Units

Part 2

Cost per Equivalent Unit

Total Materials Cost = Beginning Materials Cost + Materials Cost Added During the Year

Total Material Costs = 10,700 + 103,700

Total Material Costs = $ 114,400

Material Cost per Equivalent Unit for Materials = 114,400 / 5,375= $ 21.284 per Unit

Total Conversion Costs = Beginning Conversion Costs + Conversion Costs Added During the Year

Total Conversion Costs = 13,900 + 180,900

Total Conversion Costs = $ 194,800

Conversion Costs per Equivalent Unit for Conversion = 194,800 / 4,530 Units

= $ 43.002 Per Unit

Cost per Unit = 21.284 + 43.002

Cost per Unit = $ 64.286

Part C

Cost of Units Transferred Out = 4,400 Units * Costs per Unit

Cost of Units Transferred Out = 4,400 * 64.286

Cost of Units Transferred Out = $ 282,858.40

Part C

Cost of Ending Work in Process in First Department

For Equivalent Units with respect to Materials = 975 Units * $ 21.284 per Unit

= $ 20,751.90

For Equivalent Units with respect to Conversion = 130 Units * $ 43.002 per Unit

= $ 5,590.26

Total Cost of Ending Work in Process Inventory = 20,751.90 + 5,590.26 = $ 26,342.16

Part E

Total Cost to Accounts for = Total Materials Cost + Total Conversion Costs

Total Cost to Accounts for = 194,800 + 114,400

Total Cost to Accounts for = $ 309,200